solve the system of equations:
8x-y+z=-23
4x+2y-3z=-16
x-3y+2z=-2

Answer by stanbon(75887)

solve the system of equations:
8x-y+z=-23
4x+2y-3z=-16
x-3y+2z=-2
--------
I used a matrix method on a TI-84 to get:
x = -3
y = 1
z = 2
============================
Cheers,
Stan H.
============